OpenStudy (anonymous):

help me please: cos(arcsin(3/5)+arctan(4)) So I have checked and doublechecked, tripple checked my work, and its still coming out wrong...

jimthompson5910 (jim_thompson5910):

to evaluate cos(arcsin(3/5)+arctan(4)), we will use the identity cos(x+y) = cos(x)*cos(y) - sin(x)*sin(y) in this case x = arcsin(3/5) y = arctan(4)

jimthompson5910 (jim_thompson5910):

did you start out that way?

OpenStudy (anonymous):

nope... i thought that once i found out what the cos was i can just add it

OpenStudy (anonymous):

let me do it your way

jimthompson5910 (jim_thompson5910):

give it a shot and post what you get

OpenStudy (anonymous):

yay i got it right!!

OpenStudy (anonymous):

i wasnt using the formula thank u

OpenStudy (anonymous):

so now that i have u can i ask u something else?

jimthompson5910 (jim_thompson5910):

sure, go ahead

OpenStudy (anonymous):

sin(pi/18) i have no idea how to solve this

OpenStudy (anonymous):

it must remain in fraction form

jimthompson5910 (jim_thompson5910):

ok one moment

jimthompson5910 (jim_thompson5910):

Any ideas satellite? I'm trying to look for a formula dealing with sin(k*pi/6) where k = 1/3, but all I can find so far are instances where k is a whole number. I'll keep looking though.

OpenStudy (anonymous):

yeah good luck with this program i tried to solve \[\frac{a}{6}+\frac{b}{3}=\frac{1}{18}\] but the problem with that is you are looking for integer solutions of \[3a+6b=1\] and that is not going to work

OpenStudy (anonymous):

i think you can do it by solving some sort of cubic equation but that is not going to be easy in this case it is easier to work in degrees, and you are looking for \[\sin(10)\] a quick google search returns this result https://answers.yahoo.com/question/index?qid=20080511045049AAOrNdH

jimthompson5910 (jim_thompson5910):

seems cruel to have you solve a cubic

OpenStudy (anonymous):

if you just type it in to wolfram you do not get the usual nested radicals etc, to make it seem like a simple application of a addition angle formula instead you get it is the root of \( 8 x^3-6 x+1 \)

OpenStudy (anonymous):

sure it is \[\sin(\frac{\pi}{18})\]?

OpenStudy (anonymous):

i did and i didnt understand how they resolved it.. sorr i meant sinpi/12

OpenStudy (anonymous):

oooooooooooooooooooooooooooooooooooooh

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

i didnt notice i had typed it out incorrectly

OpenStudy (anonymous):

\[\frac{\pi}{12}\] is half of \(\frac{\pi}{6}\) use the "half angle" formula

OpenStudy (anonymous):

yeah i am trying to do that right now but i am not sure how i would get cosA but let me see

jimthompson5910 (jim_thompson5910):

you dont need to find cos(pi/12) instead, find cos(pi/6)

OpenStudy (anonymous):

all you need is the cosine of \(\frac{\pi}{6}\) which is not hard to find

OpenStudy (anonymous):

ohhhh i got it

jimthompson5910 (jim_thompson5910):

since the formula is \[\Large \sin\left(\frac{A}{2}\right) = \sqrt{\frac{1-\cos(A)}{2}}\] some books state it as \[\Large \sin(A) = \sqrt{\frac{1-\cos(2A)}{2}}\]

OpenStudy (anonymous):

ok so simple basic question here.. i insert the numbers... and i get \[\frac{ \sqrt{1-(\sqrt{3}/2)} }{ 2 }\]

jimthompson5910 (jim_thompson5910):

yes, now simplify

OpenStudy (anonymous):

wouldnt there be a sq rt over the denominator?

OpenStudy (anonymous):

i mean on the two

jimthompson5910 (jim_thompson5910):

oh yeah it should be, just noticed that

OpenStudy (anonymous):

but according to wolfram that is the answer

jimthompson5910 (jim_thompson5910):

hmm let me see what I get

OpenStudy (anonymous):

no sorry according to wolfram its sqrt2-sqrt3 / 2

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

ok this is what I get (it's a lot, but hopefully it all makes sense) \[\Large \sin\left(\frac{\pi}{12}\right) = \sin\left(\frac{1}{2}*\frac{\pi}{6}\right)\] \[\Large \sin\left(\frac{1}{2}*A\right) = \sqrt{\frac{1-\cos(A)}{2}}\] \[\Large \sin\left(\frac{1}{2}*\frac{\pi}{6}\right) = \sqrt{\frac{1-\cos\left(\frac{\pi}{6}\right)}{2}}\] \[\Large \sin\left(\frac{1}{2}*\frac{\pi}{6}\right) = \sqrt{\frac{1-\frac{\sqrt{3}}{2}}{2}}\] \[\Large \sin\left(\frac{1}{2}*\frac{\pi}{6}\right) = \sqrt{\frac{2}{2}*\frac{1-\frac{\sqrt{3}}{2}}{2}}\] \[\Large \sin\left(\frac{1}{2}*\frac{\pi}{6}\right) = \sqrt{\frac{2*\left(1-\frac{\sqrt{3}}{2}\right)}{2*2}}\] \[\Large \sin\left(\frac{1}{2}*\frac{\pi}{6}\right) = \sqrt{\frac{2*\left(1\right)-2*\left(\frac{\sqrt{3}}{2}\right)}{2*2}}\] \[\Large \sin\left(\frac{1}{2}*\frac{\pi}{6}\right) = \sqrt{\frac{2-\sqrt{3}}{4}}\] \[\Large \sin\left(\frac{1}{2}*\frac{\pi}{6}\right) = \frac{\sqrt{2-\sqrt{3}}}{\sqrt{4}}\] \[\Large \sin\left(\frac{1}{2}*\frac{\pi}{6}\right) = \frac{\sqrt{2-\sqrt{3}}}{2}\] \[\Large \sin\left(\frac{\pi}{12}\right) = \frac{\sqrt{2-\sqrt{3}}}{2}\]

OpenStudy (anonymous):

wait why are u multiplying the two?

jimthompson5910 (jim_thompson5910):

with this step here? \[\Large \Large \sin\left(\frac{1}{2}*\frac{\pi}{6}\right) = \sqrt{\frac{2}{2}*\frac{1-\frac{\sqrt{3}}{2}}{2}}\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

I multiplied top and bottom of the fraction by the LCD 2 to clear out that inner fraction

jimthompson5910 (jim_thompson5910):

it's the same as multiplying by 1 since 2/2 = 1

OpenStudy (anonymous):

ok got it makes more sense... thank u so much

OpenStudy (anonymous):

yeah i know... i just didnt see it like that

jimthompson5910 (jim_thompson5910):

glad its clicking now

OpenStudy (anonymous):

thank u so so much

jimthompson5910 (jim_thompson5910):

you're welcome